OpenStudy (lgbasallote):

\(\Huge \color{darkblue}{\mathfrak{LGBARIDDLE}}\) Basic arithmetic \(\ddot \smile\) 5 - 2 = 3 6 - 2 = 4 7 - 4 = 3 5 - 1 = 3 8 - 3 = ? Explain

OpenStudy (karatechopper):

8-3=5 u have 8 cookies, u ate 3 of them. u count how many are left, u see that there are 5 left.

OpenStudy (anonymous):

\[\Huge \color{darkblue}{\mathfrak{what}}\] \[\Huge \color{darkred}{\mathfrak{in }}\] \[\Huge \color{darkgreen}{\mathfrak{the }}\] \[\Huge \color{yellow}{\mathfrak{world?}}\]

OpenStudy (lgbasallote):

\(\Huge \color{lightblue}{4} \color{pink}{2}\)

OpenStudy (karatechopper):

what the 42 for?

OpenStudy (lgbasallote):

42 -> answer to life lol

OpenStudy (karatechopper):

???

OpenStudy (zepp):

Not again?! D:

OpenStudy (anonymous):

42 is the answer to life, the universe, and everything.

OpenStudy (lgbasallote):

haha @zepp dont you just love arithmetic :P

OpenStudy (zepp):

4?

OpenStudy (lgbasallote):

how'd you get it?

OpenStudy (zepp):

EZ :D

OpenStudy (karatechopper):

???????????????????????????????????????????????

OpenStudy (zepp):

Five(4) - Two(3) = 1(ONE) 3 Letters Six(3) - Two(3) = 0(ZERO) 4 Letters Seven(5) - Four(4) = 1(ONE) 3 Letters Five(4) - One(3) = 1(ONE) 3 Letters Eight(5) - Three(5) = 0(ZERO) 4 Letters You shall not fry my brain :D

OpenStudy (lgbasallote):

hah guess not today =))

OpenStudy (zepp):

haha ;P Really impressive, creating riddles like that :O

OpenStudy (lgbasallote):

i like arithmetic \(\ddot \smile\)